OpenStudy (anonymous):

Find a polynomial function with the given zeros -1,7, - radical 2, and radical 2

OpenStudy (mathmale):

You are given FOUR zeros: {-1,7,-Sqrt(2) and Sqrt(2)}. Each one corresponds to a factor of your polynomial. for example, -1 corresponds to the factor (x-[-1]), or (x+1). what are the other three factors? Multiply all four factors together to obtain the polynomial in question.

OpenStudy (anonymous):

Would the answer be: (X+1)(x-7)(x-sqrt(2))(x+sqrt(2)) X^2-6x-7)(x^2-sqrt(2))?

OpenStudy (mathmale):

Nancyo: Please note that (x-sqrt(2))(x+sqrt(2)) = x^2-2. Try again, OK?

OpenStudy (mathmale):

\[(a-b)(a+b)=a^2-b^2.\]

OpenStudy (mathmale):

\[(x-\sqrt{2})(x+\sqrt{2})=?\]

OpenStudy (anonymous):

X^2-2i?

OpenStudy (mathmale):

Building upon what you already have: (X+1)(x-7)(x-sqrt(2))(x+sqrt(2))\[(x+1)(x-7)(x^2-2)=?\]

OpenStudy (anonymous):

X^4-7x^2-14?

OpenStudy (anonymous):

X^4-6x^3-9x^2-12x+14?